OpenStudy (anonymous):

Find the solution to the equation 64(x – 3) = 42x. explain the method used to solve this equation.

OpenStudy (callisto):

1. expand the terms 64(x – 3) = 42x. 64x - 3(64) =42x 64x - 192 = 42x 2. Subtract 64x from both sides 3. divide both sides by coefficient of x Can you do it?

OpenStudy (anonymous):

i think i did it wrong i got a decimal ..

OpenStudy (callisto):

it is a not an integer

OpenStudy (anonymous):

i'm lost .

OpenStudy (callisto):

where?

OpenStudy (anonymous):

9/5?

OpenStudy (callisto):

nope

OpenStudy (anonymous):

well i have no idea what i'm doing then .

OpenStudy (callisto):

64x - 192 = 42x Subtract 64x from both sides => what do you get?